What is poission bracket


1
Expert's answer
2021-02-15T17:38:48-0500

Answer

In fact, the Poisson bracket can be defined for any two functions defined in phase space:

"[f, g]=\Sigma(\frac{df}{dp}\frac{dg}{dq}-\frac{df}{dq}\frac{dg}{dp})"

For canonical transformation

"[f, g]=1"